higherorder patterns? ? ? ? ? ? ? ? ? ? ? λxEx E λxE λxM M x λxM M does x λ strictness invariance linear patterns λ ? ?? ?? ?? l ?? ?? ?? ?? ?? ? ? ?? ? uncover ? ?? L t

L(x) = ∅
L(f())=
{g() | g ∈ Σf} ∪
{f(z_{1}z_{i−1}, s, z_{i+1}, …, z_{n}) | s ∈_{L} (t_{i}), i ∈ [1, n]}

disunification ? ? l l ? M ∈ Λ_{I} x ∈ F V (M) M ∈ Λ_{I} Λ_{I} Λ K Λ Λ_{I} K l I ? 2.2.2− 2.2.5 9 ? K I, S, B, C ? ? only

Λ^{k} k Λ^{1} = Λ^{u}∩ Λ_{I} ? 1

Some early work in the twenties in the Soviet Union was, at the time, not accessible.